# Predictive Assessment and Comparison of Bayesian Survival Models for Cancer Recurrence

### v1

@@ Abstract
Complex data features, such as unmodelled censored event times and variables with time-dependent effects, are common in cancer recurrence studies and pose challenges for Bayesian survival modelling. However, current methodologies for predictive model checking and comparison often fail to adequately address these features. This paper bridges that gap by introducing new, targeted recommendations for predictive assessment and comparison of Bayesian survival models for cancer recurrence. Our recommendations cover a variety of different scenarios and models. Accompanying code together with our implementations to open source software help in replicating the results and applying our recommendations in practice.
@@
Bayesian inference; survival analysis; predictive model checking; predictive model comparison; cross-validation; cancer recurrence.
@@ 1 Introduction
The motivation for this work is to improve the predictive model checking and the predictive model comparison of Bayesian survival models with the aim of analysing the gastrointestinal stromal tumour (GIST) data set that was collected over multiple decades in a collaborative effort of multiple research centres around the world. The data set has interesting aspects related to predictive model checking and predictive model comparison that are common in cancer studies but have not received much attention in the Bayesian literature and software implementations. A large portion of the event times in the GIST data set are censored, but the independent censoring process is not known nor modelled. This means that the full generative model is not available, which complicates predictive model checking and predictive model comparison. The GIST data set also has variables with clear time-dependent effects, which encourages the building of more complex models that further complicate predictive model comparison.
All GIST-related analysis presented in this paper is performed on simulated data, since we are prohibited from publishing results from the real data set. In our other examples, we use simulated data or the Rotterdam breast cancer data set (royston2013). All the code and data used in this paper are freely available.  [^1] Code and data required to replicate the results in this paper are freely available at https://github.com/Sakuski/Bayesian-Cancer-Recurrence-Prediction.
@@ 1.1 Our Contributions
We present new recommendations for predictive model checking and predictive model comparison in Bayesian survival modelling. We give recommendations on how to perform predictive checking for different survival models in case there are no censored event times in the data, and in case there are censored event times in the data and the censoring process is not modelled. We also give recommendations for performing predictive model comparison when some of the models cause problems in the regular Pareto Smoothed Importance Sampling Leave-One-Out Cross-Validation (PSIS-LOO CV) based comparison. The recommendations are summarised in Section 6 and justified in Section 3 and Section 4. Some of our recommendations were included as software implementations in the release v.1.13.0 of bayesplot (bayesplot).
@@ 1.2 Relation to Previous Work
Although the roots of survival analysis go back to the 17th century (graunt1662), it is the ground-breaking work by kaplan1958, cox1972, nelson1972, and aalen1978 in the later part of the 20th century that leads to the modern framework. In the early years, survival analysis mainly followed the frequentist approach, but as computational capabilities grew larger, the Bayesian approach became a common alternative. Notable early contributions in this area include the work by kalbfleisch1978 and ibrahim2001.
In recent years, the concept of Bayesian workflow (gelman2020_workflow) as a structural process has been established to help with practical aspects of Bayesian analysis. Different parts of the Bayesian workflow, including inference, predictive model checking, and predictive model comparison, have been automated in software tools, such as brms (brms2017; brms2018; brms2021), bayesplot (bayesplot; gabry2019), and loo (vehtari2024_loo), respectively.
Bayesian methodology has been applied in previous work related to gastrointestinal stromal tumours. For example, joensuu2012 utilised a modified Cox proportional hazards model fitted with Bayesian inference to non-linearly model recurrence free survival. In addition, joensuu2014 modelled the hazard of recurrence of GIST with Bayesian methods to optimise the timing of CT scans.
The relationship of our work and previous research can be summarised in the following words. We take a close look at two important steps in the established Bayesian workflow: predictive model checking and predictive model comparison. We improve these steps in the context of survival modelling through our contributions. We also improve and give improvement suggestions to the relevant software tools in this area. In addition, we continue to apply Bayesian survival modelling in the analysis of data on gastrointestinal stromal tumours.
@@ 2 Survival Models
In this paper, we focus only on models that are supported by the brms package (brms2017; brms2018; brms2021). The brms framework allows for the inclusion of censoring and left-truncation in survival modelling. The possible model types include the common proportional hazards (PH) and accelerated failure time (AFT) models, but also a Bernoulli model that requires a bit more manual crafting.
@@
For each individual and for each discrete time interval in which they are observed and at risk, we model the probability of cancer recurrence using a Bernoulli distribution. This means that within any given interval, the outcome for a subject is binary: either a recurrence occurs, or it does not. A key point is that time intervals after a subject has experienced the event (recurrence) or has been censored (e.g., died from other causes, lost to follow-up) are excluded from the analysis.
@@
This modelling approach requires the data set to be transferred to a long format. Instead of having one row per subject, we create a data point for each discrete time interval for each individual. In this format, a single subject is represented by multiple rows, one for each time interval in which they were observed and at risk. Time-invariant covariates remain constant across all rows for the same subject. The main advantage of the long data format is its ability to incorporate time-dependent covariates, which would be impossible in a standard short data format.
@@ 3 Predictive Checking of Survival Models
If the generative model is available, it can be used to perform predictive model checks. Predictive model checking is an important part of the Bayesian workflow. Before fitting the model, it is useful to simulate data from the prior predictive distribution to check that the priors combined with the likelihood produce reasonable data. This is called prior predictive checking (gelman2013). Similarly, after fitting the model, it is important to verify that the observed data are plausible under the fitted model. This is referred to as posterior predictive checking (gelman2013). Posterior predictive checking can detect model misspecification and prevent problematic scenarios in leave-one-out cross-validation (sivula2020). Predictive model checks can be performed visually (gabry2019), which is the focus of this section.
@@ 3.1 Predictive Checking of Parametric Survival Models in the Absence of Censored Event Times
Figure 3 shows a more complicated PIT-ECDF plot, which is based on the following mathematical result: Assume a continuous random variable $X$ for which the cumulative distribution function (CDF) is $F_{X}$. Then, the transformed random variable $Y=F_{X}(X)$ follows a standard uniform distribution (i.e., $Y\sim U(0,1)$). In practice, a PIT value is calculated for each observation $y_{i}$ as a proportion of its corresponding predictive draws $y_{i}^{\text{rep}}$ that are less than or equal to the observed value. Formally:
@@
where $S$ is the number of predictive draws and $I(\cdot)$ is an indicator function. After the PIT values are calculated, their empirical cumulative distribution function (ECDF) is plotted. If the model is well calibrated, the ECDF of the PIT values should be close to a straight line from 0 to 1, as this indicates that the PIT values follow a standard uniform distribution.
@@
Figure 4 displays the Kaplan-Meier overlay plot. The plot shows the Kaplan-Meier estimator (kaplan1958) of the survival function $S(t)$ for the observations. The estimator is defined as
@@
Figure 4: The Kaplan-Meier overlay plot shows the Kaplan-Meier survival curve (kaplan1958) of the observations with a dark colour and the empirical CCDF estimates of the predictive draws with a lighter colour. If the survival curve deviates from the empirical CCDF estimates of the predictive draws, there is probably some lack of fit.
There is a small detail in Kaplan-Meier overlay plots that has received less attention in the software implementations. When the data set contains delayed entries, they should be taken into account in the Kaplan-Meier overlay plot. Individuals who enter the study late are conditioned on the fact that they have not experienced the event before their entry time point, which means that they do not represent the population at the beginning of the study. Therefore, these individuals cannot be considered at risk before entering the study. That is, $n_{i}$, which represents the number of individuals at risk at the time point $t_{i}$ in Equation 3, should not count individuals who have entered the study after the time point $t_{i}$. Otherwise, $n_{i}$ would be too high, which, as seen in the equation, would result in overestimated survival probabilities at early time points. The impact of this mistake is illustrated in Figure 5.
@@ 3.2 Predictive Checking of Parametric Survival Models in the Presence of Censored Event Times
When the censoring process is not modelled, some diagnostic plots become useless. Examples of these types of plot are the intervals plot and the PIT-ECDF plot, which are shown in Figure 6(b). These two diagnostic plots always indicate a lack of fit when there are censored event times in the data and the censoring process is not modelled. Therefore, these diagnostic plots do not provide much value as a predictive model check in this kind of case, and we recommend not using them unless the censored event times are dealt with.
@@
Figure 8 shows an improved version of the Kaplan-Meier overlay plot in which only 20% of the extrapolation is shown. In this visualisation, the main focus is on the observed data, which is the most interesting part of the figure. The zoomed plot reveals that the Kaplan-Meier survival curve for the observations deviates from the empirical CCDF estimates for the predictive draws, which would have been difficult to notice in Figure 7.
@@ 3.3 Predictive Checking of the Bernoulli Model
In contrast to other survival models, the Bernoulli model has a binary outcome. For binary outcomes, a common mistake when performing a predictive model check is the use of bar graphs. Bar graphs are plots that show bars corresponding to frequencies of observed values and point intervals of predictive means and centre quantiles. According to sailynoja2025, bar graphs are often meaningless as a predictive model check for binary data, since even the simplest intercept only models can perfectly model the proportions of the values, and bar graphs do not allow inspection of any other aspects, such as under- and overconfidence.
Figure 10: Example of a bar graph.
To create a more comprehensive model check for binary outcomes, sailynoja2025 recommend the PAV-adjusted calibration plots applying the work by dimitriadis2021. The idea of PAV-adjusted calibration plots is to compute conditional event probabilities (CEPs) using the posterior predictive means for each observation and then to form consistency bands that show how much variation should be expected from a perfectly calibrated model. CEPs denote the probability that the event occurs given that the model has predicted some probability. In a perfectly calibrated model, the CEPs are equal to the predicted probabilities. The CEPs are calculated in the PAV-adjusted calibration using an algorithm called pool adjacent violators (PAV) (ayer1955), which the calibration plot is named after.
@@ 3.4 Predictive Model Checks Suitable for Any Survival Model
It is also possible to create predictive model checks that are suitable for any survival model in the presence and absence of censored event times regardless of the target variable of the model. However, these types of predictive model checks require compromises that can potentially lose some information.
@@ 4 Predictive Comparison of Survival Models
To compare predictive performance between models, a common approach is to calculate predictive densities using log scores (geisser1979; bernardo1994; gneiting2007) as a utility. Furthermore, vehtari2017 define expected log-predictive densities as a measure of predictive performance for a new data set of $n$ observations taken one at a time
@@
The state-of-the-art method for computing estimates for $\mathrm{elpd_{LOO}}$ (vehtari2017) uses pareto-smoothed importance sampling (PSIS) (vehtari2024_psis). In practice, the PSIS-LOO CV based model comparison can be performed using the functions in the loo R package (vehtari2024_loo). The loo() function can be used to calculate the elpd estimate $\widehat{\mathrm{elpd}}$ for each model. The loo_compare() function can then compare the elpd estimates between models. We denote the differences in the elpd estimates between models by $\Delta\widehat{\mathrm{elpd}}$. For the best model, $\Delta\widehat{\mathrm{elpd}}=0$, and for the other models, $\Delta\widehat{\mathrm{elpd}}$ closer to zero is better. loo_compare() also reports the standard error related to the comparison. According to vehtari2017, the standard error of $\mathrm{\widehat{elpd}}$ is defined as
@@
where $n$ is the number of data points. As justified by sivula2020, the standard error can be used to quantify the uncertainty in the comparison of models. The standard error related to the comparison of models A and B is
@@
If $\Delta\widehat{\mathrm{elpd}}$ is less than two standard errors from zero, the model is considered indistinguishable in predictive performance compared to the best model (riha2024).
@@ 4.2 Predictive Comparison of PH and AFT Models in the Presence of Censored Event Times
When the data set contains censored event times, the predictive model comparison becomes more difficult. In case the data point $y_{i}$ is censored, $p_{k}(y_{i}\mid\theta)$ in Equation 5 is interpreted in terms of the true event time $y_{i}^{\ast}$ as
1. $p_{k}(y_{i}^{\ast}>a\mid\theta)$, if $y_{i}$ is right-censored at $a$;
2. $p_{k}(y_{i}^{\ast}<b\mid\theta)$, if $y_{i}$ is left-censored at $b$; and
3. $p_{k}(a<y_{i}^{\ast}<b\mid\theta)$, if $y_{i}$ is interval-censored with endpoints $a$ and $b$.
This means that pointwise elpds for censored data points are not densities, but probabilities. The phenomenon can be clearly illustrated when calculating pointwise elpds for two models, which are otherwise the same, but the time scales of the event times are different. Now, pointwise elpds stay the same for the censored data points but change for the observed ones.
To create an example, we fitted a Weibull AFT model on a data set where the event time is in days, and a significant portion of the event times are right-censored. After that, we calculated pointwise elpds and plotted their histogram in Figure 15(a). There are two clear clusters in the figure. The right cluster contains the pointwise elpds for the censored data points, and the left cluster has the pointwise elpds for the non-censored data points.
@@ 2. $p_{k}(y_{i}^{\ast}<b\mid\theta)$, if $y_{i}$ is left-censored at $b$; and
It is not possible to convert probabilities into densities, but it is possible to convert densities into probabilities. The conversion can be performed by integrating the densities over different time intervals. The shorter the time intervals are, the less information is lost. Once the densities have been converted into probabilities, the PSIS-LOO CV can be applied without issues.

### v2

@@ Abstract
Complex data features, such as unmodelled censored event times and variables with time-dependent effects, are common in cancer recurrence studies and pose challenges for Bayesian survival modelling. Current methodologies for predictive model checking and comparison often fail to adequately address these features. This paper bridges that gap by introducing new, targeted recommendations for predictive assessment and comparison of Bayesian survival models. Our recommendations cover a variety of different scenarios and models. Accompanying code together with our implementations to open source software help in replicating the results and applying our recommendations in practice.
@@
Bayesian inference; survival analysis; predictive model checking; predictive model comparison; cross-validation
@@ 1 Introduction
The motivation for this work is to improve the predictive model checking and the predictive model comparison of Bayesian survival models with the aim of analysing the gastrointestinal stromal tumour (GIST) data set that was collected over multiple decades in a collaborative effort of multiple research centres around the world. The data set has interesting aspects related to predictive model checking and predictive model comparison that are common in survival analysis but have not received much attention in the Bayesian literature and software implementations. A large portion of the event times in the GIST data set are censored, but the independent censoring process is not known nor modelled. This means that the full generative model is not available, which complicates predictive model checking and predictive model comparison. The GIST data set also has variables with clear time-dependent effects, which encourages the building of more complex models that further complicate predictive model comparison.
All GIST-related analysis presented in this paper is performed on simulated data, since we are prohibited from publishing results from the real data set. In our other examples, we use simulated data or the Rotterdam breast cancer data set (Royston and Altman, 2013). All the code and data used in this paper are freely available.  [^1] Code and data required to replicate the results in this paper are freely available at https://github.com/Sakuski/Bayesian-Cancer-Recurrence-Prediction.
@@ 1.1 Our contributions
We present new recommendations for predictive model checking and predictive model comparison in Bayesian survival modelling. We give recommendations on how to perform predictive checking for different survival models in case there are no censored event times in the data, and in case there are censored event times in the data and the censoring process is not modelled. We also give recommendations for performing predictive model comparison when some of the models cause problems in the regular Pareto smoothed importance sampling leave-one-out cross-validation (PSIS-LOO CV) based comparison. The recommendations are summarised in Section 6 and justified in Section 3 and Section 4. Some of our recommendations were included as software implementations in the release v.1.13.0 of bayesplot R package (Gabry and Mahr, 2025).
@@ 1.2 Relation to previous work
Although the roots of survival analysis go back to the 17th century (Graunt, 1662), it is the ground-breaking work by Kaplan and Meier (1958), Cox (1972), Nelson (1972), and Aalen (1978) in the later part of the 20th century that leads to the modern framework. In the early years, survival analysis mainly followed the frequentist approach, but as computational capabilities grew larger, the Bayesian approach became a common alternative. Notable early contributions in this area include the work by Kalbfleisch (1978) and Ibrahim et al. (2001).
In recent years, the concept of Bayesian workflow (e.g., Gelman et al., 2020) as a structural process has been established to help with practical aspects of Bayesian analysis. Different parts of the Bayesian workflow, including inference, predictive model checking, and predictive model comparison, have been automated in software packages, such as brms (Bürkner, 2017, 2018, 2021), bayesplot (Gabry and Mahr, 2025; Gabry et al., 2019), and loo (Vehtari et al., 2024b), respectively.
Bayesian methodology has been applied in previous work related to gastrointestinal stromal tumours. For example, Joensuu et al. (2012) utilised a modified Cox proportional hazards model fitted with Bayesian inference to non-linearly model recurrence free survival. In addition, Joensuu et al. (2014) modelled the hazard of recurrence of GIST with Bayesian methods to optimise the timing of CT scans.
The relationship of our work and previous research can be summarised in the following words. We take a close look at two important steps in the established Bayesian workflow: predictive model checking and predictive model comparison. We improve these steps in the context of survival modelling through our contributions. We also improve and give improvement suggestions to the relevant software tools in this area. In addition, we continue to apply Bayesian survival modelling in the analysis of simulated data on recurrence of gastrointestinal stromal tumours.
@@ 2 Survival models
In this paper, we focus only on models that are supported by the brms R package (Bürkner, 2017, 2018, 2021). The brms framework allows for the inclusion of censoring and left-truncation in survival modelling. The possible model types include the common proportional hazards (PH) and accelerated failure time (AFT) models, but also a Bernoulli model that requires a bit more manual crafting.
@@
For each individual and for each discrete time interval in which they are observed and at risk, we model the probability of cancer recurrence using a Bernoulli model. This means that within any given interval, the outcome for a subject is binary: either a recurrence occurs, or it does not. A key point is that time intervals after a subject has experienced the event (recurrence) or has been censored (e.g., died from other causes, lost to follow-up) are excluded from the analysis.
@@
This modelling approach requires the data set to be transferred to a long format. Instead of having one row per subject, we create a data point for each discrete time interval for each individual. In this format, a single subject is represented by multiple rows, one for each time interval in which they were observed and at risk. Time-invariant covariates remain constant across all rows for the same subject. The main advantage of the long data format is its ability to easily incorporate time-dependent covariates.
@@ 3 Predictive checking of survival models
If the generative model is available, it can be used to perform predictive model checks. Predictive model checking is an important part of the Bayesian workflow. Before fitting the model, it is useful to simulate data from the prior predictive distribution to check that the priors combined with the likelihood produce reasonable data. This is called prior predictive checking (Gelman et al., 2013). Similarly, after fitting the model, it is important to verify that the observed data are plausible under the fitted model. This is referred to as posterior predictive checking (Gelman et al., 2013). Posterior predictive checking can detect model misspecification and prevent problematic scenarios in leave-one-out cross-validation (Sivula et al., 2025). Predictive model checks can be performed visually (Gabry et al., 2019), which is the focus of this section.
@@ 3.1 Predictive checking of parametric survival models in the absence of censored event times
Figure 3 shows a PIT-ECDF plot, which is based on the following mathematical result: Assume a continuous random variable $X$ for which the cumulative distribution function (CDF) is $F_{X}$. Then, the transformed random variable $Y=F_{X}(X)$ follows a standard uniform distribution (i.e., $Y\sim U(0,1)$). In practice, a PIT value is often computed for each observation $y_{i}$ as a proportion of its corresponding predictive draws $y_{i}^{\text{rep}}$ that are less than or equal to the observed value. Formally:
@@
where $S$ is the number of predictive draws and $I(\cdot)$ is an indicator function. After the PIT values are computed, their empirical cumulative distribution function (ECDF) is plotted. If the model is well calibrated, the ECDF of the PIT values should be close to a straight line from 0 to 1, as this indicates that the PIT values follow a standard uniform distribution.
@@
Figure 4 displays the Kaplan-Meier overlay plot. The plot shows the Kaplan-Meier estimator (Kaplan and Meier, 1958) of the survival function $S(t)$ for the observations. The estimator is defined as
@@
Figure 4: The Kaplan-Meier overlay plot shows the Kaplan-Meier survival curve (Kaplan and Meier, 1958) of the observations with a dark colour and the empirical CCDF estimates of the predictive draws with a lighter colour. If the survival curve deviates from the empirical CCDF estimates of the predictive draws, there is probably some lack of fit.
There is a small detail in Kaplan-Meier overlay plots that has received less attention in the software implementations. When the data set contains delayed entries, that should be taken into account in the Kaplan-Meier overlay plot. Individuals who enter the study late are conditioned on the fact that they have not experienced the event before their entry time point, which means that they do not represent the population at the beginning of the study. Therefore, these individuals cannot be considered at risk before entering the study. That is, $n_{i}$, which represents the number of individuals at risk at the time point $t_{i}$ in Equation 3, should not count individuals who have entered the study after the time point $t_{i}$. Otherwise, $n_{i}$ would be too high, which, as seen in the equation, would result in overestimated survival probabilities at early time points. The impact of this mistake is illustrated in Figure 5.
@@ 3.2 Predictive checking of parametric survival models in the presence of censored event times
When the censoring process is not modelled, some diagnostic plots become useless. Examples of these types of plot are the intervals plot and the PIT-ECDF plot, which are shown in Figure 6. These two diagnostic plots always indicate a lack of fit when there are censored event times in the data and the censoring process is not modelled. Therefore, these diagnostic plots do not provide much value as a predictive model check in this kind of case, and we recommend not using them unless the censored event times are dealt with.
@@
Figure 8 shows an improved version of the Kaplan-Meier overlay plot in which the extrapolation is shown 20% beyond the furthest observed time point. In this visualisation, the main focus is on the observed data, which is the most interesting part of the figure. The zoomed plot reveals that the Kaplan-Meier survival curve for the observations deviates from the empirical CCDF estimates for the predictive draws, which would have been difficult to notice in Figure 7.
@@ 3.3 Predictive checking of the Bernoulli model
In contrast to other survival models, the Bernoulli model has a binary outcome. For binary outcomes, a common mistake when performing a predictive model check is the use of bar graphs. Bar graphs are plots that show bars corresponding to frequencies of observed values and point intervals of predictive means and centre quantiles. According to Säilynoja et al. (2025), bar graphs are often meaningless as a predictive model check for binary data, since even the simplest intercept only models can perfectly model the proportions of the values, and bar graphs do not allow inspection of any other aspects, such as under- and overconfidence.
Figure 10: Example of a bar graph for binary target.
To create a more comprehensive model check for binary outcomes, Säilynoja et al. (2025) recommend the PAV-adjusted calibration plots applying the work by Dimitriadis et al. (2021). The idea of PAV-adjusted calibration plots is to compute conditional event probabilities (CEPs) using the posterior predictive means for each observation and then to form consistency bands that show how much variation should be expected from a perfectly calibrated model. CEPs denote the probability that the event occurs given that the model has predicted some probability. In a perfectly calibrated model, the CEPs are equal to the predicted probabilities. The CEPs are computed in the PAV-adjusted calibration using an algorithm called pool adjacent violators (PAV) (Ayer et al., 1955), which the calibration plot is named after.
@@ 3.4 Predictive model checks suitable for any survival model
It is also possible to create predictive model checks that are suitable for any survival model in the presence and absence of censored event times regardless of the target variable of the model. These types of predictive model checks require compromises that can potentially lose some information.
@@ 4 Predictive comparison of survival models
To compare predictive performance between models, a common approach is to compute predictive densities using log scores (Geisser and Eddy, 1979; Bernardo and Smith, 1994; Gneiting and Raftery, 2007) as a utility. Furthermore, Vehtari et al. (2017) define expected log-predictive densities as a measure of predictive performance for a new data set of $n$ observations taken one at a time
@@
The state-of-the-art method for computing estimates for $\mathrm{elpd_{LOO}}$ (Vehtari et al., 2017) uses Pareto smoothed importance sampling (PSIS) (Vehtari et al., 2024a). In practice, the PSIS-LOO CV based model comparison can be performed using the functions in the loo R package (Vehtari et al., 2024b). The loo() function can be used to compute the elpd estimate $\widehat{\mathrm{elpd}}$ for each model. The loo_compare() function can then compare the elpd estimates between models. We denote the differences in the elpd estimates between models by $\Delta\widehat{\mathrm{elpd}}$. For the best model, $\Delta\widehat{\mathrm{elpd}}=0$, and for the other models, $\Delta\widehat{\mathrm{elpd}}$ closer to zero is better. loo_compare() also reports the standard error related to the comparison. According to Vehtari et al. (2017), the standard error of $\mathrm{\widehat{elpd}}$ is defined as
@@
where $n$ is the number of data points. As justified by Sivula et al. (2025), the standard error can be used to quantify the uncertainty in the comparison of models. The standard error related to the comparison of models A and B is
@@
If $\Delta\widehat{\mathrm{elpd}}$ is less than two standard errors from zero, the model is considered indistinguishable in predictive performance compared to the best model (Riha et al., 2024).
@@ 4.2 Predictive comparison of PH and AFT models in the presence of censored event times
When the data set contains censored event times, the predictive model comparison becomes more complicated. In case the data point $y_{i}$ is censored, $p_{k}(y_{i}\mid\theta)$ in Equation 5 is interpreted in terms of the true event time $y_{i}^{\ast}$ as
  1. 1.
    $p_{k}(y_{i}^{\ast}>a\mid\theta)$, if $y_{i}$ is right-censored at $a$;
  2. 2.
    $p_{k}(y_{i}^{\ast}<b\mid\theta)$, if $y_{i}$ is left-censored at $b$; and
  3. 3.
    $p_{k}(a<y_{i}^{\ast}<b\mid\theta)$, if $y_{i}$ is interval-censored with endpoints $a$ and $b$.
This means that pointwise elpds for censored data points are not densities, but probabilities. The phenomenon can be clearly illustrated when computing pointwise elpds for two models, which are otherwise the same, but the time scales of the event times are different. Now, pointwise elpds stay the same for the censored data points but change for the observed ones.
To create an example, we fitted a Weibull AFT model on a data set where the event time is in days, and a significant portion of the event times are right-censored. After that, we computed pointwise elpds and plotted their histogram in Figure 15(a). There are two clear clusters in the figure. The right cluster contains the pointwise elpds for the censored data points, and the left cluster has the pointwise elpds for the non-censored data points.
@@
It is not possible to convert probabilities into densities, but it is possible to convert densities into probabilities. Once the densities have been converted into probabilities, the PSIS-LOO CV can be applied without issues. The conversion can be performed by integrating the densities over different time intervals. In general, the shorter the time intervals are, the less information is lost. However, when choosing the interval length in practice, one must evaluate the tradeoff between preserving information and keeping the computational cost manageable, since shorter intervals also mean higher computational cost. A good starting point for deciding the interval length is to consider the data measurement resolution, as intervals shorter than the smallest possible time increment in the data do not add additional value compared to intervals that match the smallest possible time increment. For example, if a new data point is created every month, it is not useful to discretise the time into days. If using intervals that match the smallest possible time increment comes with overwhelming computational cost, one can gradually increase the interval length bearing in mind that each increase potentially loses some information. In order to consider the interval length to be information preserving, one should be able to assume that the probability density function remains approximately constant within each interval.